# Algorithms for Optimizing Acyclic Queries

### v1

@@ 6 Converting a Binary Join Plan to a Join Tree
To prove the “only if” direction by contradiction, we assume that a connected left-deep linear plan for a $\gamma$-acyclic query is not the reverse of a GYO-reduction order. By proposition˜38, there is an orphan $\mathring{r}$ in the plan. Assuming that $\mathring{r}$ is the $t$-th relation in the plan, we now consider the prefix of the plan up to $\mathring{r}$, namely $\mathsf{pre}({t})=r_{1},\ldots,r_{t-1},\mathring{r}$. Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, lemma˜40 guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$. There are at least two dangling variables $\bar{x}\in\bar{r}\cap\mathring{r}\setminus\hat{r}$ and $\hat{x}\in\hat{r}\cap\mathring{r}\setminus\bar{r}$.
@@ 7 Conclusion and Future Work
Our results also raise theoretical questions. Can join tree enumeration achieve worst-case constant delay? We proved Berge-acyclicity sufficient for the existence and uniqueness of the canonical join tree, but it is not necessary, and $\gamma$-acyclicity is insufficient as shown in fig.˜6 of appendix˜F. What is the precise characterization of hypergraphs that admit a unique canonical join tree for any root, or for some root?
@@ Appendix A Extended Preliminaries
Considering the hypergraph $H_{6}$ shown in fig.˜2(a), we run algorithm˜1 from the hyperedge $P$. It is labeled as $r_{1}$ and all its vertices are marked. $P$ is also tentatively assigned as the parent of five hyperedges such that $p(S)=p(T)=p(U)=p(W)=p(Y)=r_{1}$. We proceed to label the next hyperedge with the most marked vertices. At this point, each unlabeled hyperedge has one marked vertex, so we can choose any of them, say $S$, and label it as $r_{2}$.
@@
We continue the process until all hyperedges are labeled and all vertices are marked. By connecting each labeled hyperedge with its parent, we construct the join tree $T^{M}$ shown in fig.˜2(d).
Notice that algorithm˜1 can generate different join trees for the same hypergraph by breaking ties differently. However, it does not guarantee to generate all possible trees. For example, it never generates the Join Tree $T^{G}$ shown in fig.˜2(c).
@@ Appendix B Data Structures for Constant Time Operations on Graphs
Given a non-tree edge $e=(r_{i},r_{j})$ we find $\mathsf{LCA}(r_{i},r_{j})=l$ in constant time by ˜43. The depth $d=0pt(l)+1$ is likewise found in constant time. The edge $e_{1}=(l,r_{1})$, where node $r_{1}=\mathsf{LA}(r_{i},d)$ is found by a constant time look-up, is adjacent to $l$ and lies on the unique path between $r_{i}$ and $r_{j}$. The same holds for edge $e_{2}=(l,r_{2})$ where node $r_{1}=\mathsf{LA}(r_{j},d)$. ∎
@@
As we iterate over each edge $(j,\omega)$ of $L[i]$, we check if the current edge $(i,j)$ is a tree edge by performing constant-time query of $T_{r}[i]$ and $T_{r}[j]$ to see if one of them is the parent of the other. From ˜44 we know that we can find the LCA edges $\lambda(i,j)=\{e_{1},e_{2}\}$ in constant time. There’s no general solution to lookup an entry in a linked list in constant time. Instead of iterating over the linked-list representation of $L$ to find the edge weights of $e_{1}$ and $e_{2}$, we can lookup their weights from the array representation of $T_{r}$ in constant time. Let $e_{1}=(r_{1},p(r_{1}))$ and $e_{2}=(r_{2},p(r_{2}))$ with weights $\omega_{1}$ and $\omega_{2}$. Because both $e_{1}$ and $e_{2}$ are tree edges, we can find their weights by querying $T_{r}[r_{1}]=(p(r_{1}),\omega_{1})$ and $T_{r}[r_{2}]=(p(r_{2}),\omega_{2})$. Each of these operations takes constant time. When we need to delete an edge $e^{*}=(i,j)$ from $L$, we first check if $e^{*}$ is a tree edge by checking $T_{r}[i]$ and $T_{r}[j]$. Deleting $e^{*}$ from the linked list takes constant time. Sliding an edge involves updating the endpoints of the edge which takes constant time.
@@ Appendix D Missing Proofs in section˜4
lemma˜15 If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
@@
lemma˜16 Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
@@
lemma˜17 Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
@@
lemma˜19 Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
@@
2. It follows from lemma˜15 and the facts above.
@@ Appendix E Missing Proofs in section˜5
proposition˜26 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
@@
lemma˜31 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
@@
lemma˜34 Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph. Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
@@
Let $r_{0}=r$, proposition˜26 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$. By theorem˜33, the path $r_{0},\ldots,r_{k}$ is the shortest path from $r_{0}$ to $r_{k}$. The variable shared by two adjacent vertices $r_{i-1},r_{i}$ is equivalent to the edge label $\chi(r_{i-1},r_{i})$. The edge labels along the path are all distinct. Otherwise, two consecutive edges with the same label allows for shortening the path by removing either, contradicting the shortest path property. Two non-consecutive edges with the same label violate the running intersection property. ∎

### v2

@@ 6 Converting a Binary Join Plan to a Join Tree
To prove the “only if” direction by contradiction, we assume that a connected left-deep linear plan for a $\gamma$-acyclic query is not the reverse of a GYO-reduction order. By Section˜6, there is an orphan $\mathring{r}$ in the plan. Assuming that $\mathring{r}$ is the $t$-th relation in the plan, we now consider the prefix of the plan up to $\mathring{r}$, namely $\mathsf{pre}({t})=r_{1},\ldots,r_{t-1},\mathring{r}$. Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, Section˜6 guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$. There are at least two dangling variables $\bar{x}\in\bar{r}\cap\mathring{r}\setminus\hat{r}$ and $\hat{x}\in\hat{r}\cap\mathring{r}\setminus\bar{r}$.
@@ 7 Conclusion and Future Work
Our results also raise theoretical questions. Can join tree enumeration achieve worst-case constant delay? We proved Berge-acyclicity sufficient for the existence and uniqueness of the canonical join tree, but it is not necessary, and $\gamma$-acyclicity is insufficient as shown in Figure˜6 of Appendix˜F. What is the precise characterization of hypergraphs that admit a unique canonical join tree for any root, or for some root?
@@ Appendix A Extended Preliminaries
Considering the hypergraph $H_{6}$ shown in Figure˜2(a), we run Algorithm˜1 from the hyperedge $P$. It is labeled as $r_{1}$ and all its vertices are marked. $P$ is also tentatively assigned as the parent of five hyperedges such that $p(S)=p(T)=p(U)=p(W)=p(Y)=r_{1}$. We proceed to label the next hyperedge with the most marked vertices. At this point, each unlabeled hyperedge has one marked vertex, so we can choose any of them, say $S$, and label it as $r_{2}$.
@@
We continue the process until all hyperedges are labeled and all vertices are marked. By connecting each labeled hyperedge with its parent, we construct the join tree $T^{M}$ shown in Figure˜2(d).
Notice that Algorithm˜1 can generate different join trees for the same hypergraph by breaking ties differently. However, it does not guarantee to generate all possible trees. For example, it never generates the Join Tree $T^{G}$ shown in Figure˜2(c).
@@ Appendix B Data Structures for Constant Time Operations on Graphs
Given a non-tree edge $e=(r_{i},r_{j})$ we find $\mathsf{LCA}(r_{i},r_{j})=l$ in constant time by Appendix˜B. The depth $d=0pt(l)+1$ is likewise found in constant time. The edge $e_{1}=(l,r_{1})$, where node $r_{1}=\mathsf{LA}(r_{i},d)$ is found by a constant time look-up, is adjacent to $l$ and lies on the unique path between $r_{i}$ and $r_{j}$. The same holds for edge $e_{2}=(l,r_{2})$ where node $r_{1}=\mathsf{LA}(r_{j},d)$. ∎
@@
As we iterate over each edge $(j,\omega)$ of $L[i]$, we check if the current edge $(i,j)$ is a tree edge by performing constant-time query of $T_{r}[i]$ and $T_{r}[j]$ to see if one of them is the parent of the other. From Appendix˜B we know that we can find the LCA edges $\lambda(i,j)=\{e_{1},e_{2}\}$ in constant time. There’s no general solution to lookup an entry in a linked list in constant time. Instead of iterating over the linked-list representation of $L$ to find the edge weights of $e_{1}$ and $e_{2}$, we can lookup their weights from the array representation of $T_{r}$ in constant time. Let $e_{1}=(r_{1},p(r_{1}))$ and $e_{2}=(r_{2},p(r_{2}))$ with weights $\omega_{1}$ and $\omega_{2}$. Because both $e_{1}$ and $e_{2}$ are tree edges, we can find their weights by querying $T_{r}[r_{1}]=(p(r_{1}),\omega_{1})$ and $T_{r}[r_{2}]=(p(r_{2}),\omega_{2})$. Each of these operations takes constant time. When we need to delete an edge $e^{*}=(i,j)$ from $L$, we first check if $e^{*}$ is a tree edge by checking $T_{r}[i]$ and $T_{r}[j]$. Deleting $e^{*}$ from the linked list takes constant time. Sliding an edge involves updating the endpoints of the edge which takes constant time.
@@ Appendix D Missing Proofs in Section˜4
Section˜4.2 If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
@@
Section˜4.2 Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
@@
Section˜4.2 Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
@@
Section˜4.2 Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
@@
2. It follows from Section˜4.2 and the facts above.
@@ Appendix E Missing Proofs in Section˜5
Section˜5 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
@@
Section˜5.1 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
@@
Section˜5.2 Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph. Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
@@
Let $r_{0}=r$, Section˜5 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$. By Theorem˜33, the path $r_{0},\ldots,r_{k}$ is the shortest path from $r_{0}$ to $r_{k}$. The variable shared by two adjacent vertices $r_{i-1},r_{i}$ is equivalent to the edge label $\chi(r_{i-1},r_{i})$. The edge labels along the path are all distinct. Otherwise, two consecutive edges with the same label allows for shortening the path by removing either, contradicting the shortest path property. Two non-consecutive edges with the same label violate the running intersection property. ∎